OpenStudy (anonymous):

A farmer looks into the barnyard at his pigs and chickens and he counts 169 heads and 398 feet how many pigs and chickens are their? I need an equation PLEASE

OpenStudy (ash2326):

@PJohnson89 how many legs a chicken has? and a pig?

OpenStudy (anonymous):

X = Pigs Y = Chickens Total heads (assume 1 per pig and 1 per chicken) = 169 Total legs (assume 4 per pig and 2 per chicken) = 398 X + Y = 169 heads 4X + 2Y = 398 legs You now have two equations with two unknowns, ready to solve. Y = 169 - X Substitute for Y 4X + 2(169 - X) = 398 4X + 338 - 2X = 398 2X = 60 X = 30 Pigs now if you still didn't understand .. raise questions ;)

OpenStudy (anonymous):

@PJohnson89 should i be more clear ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay .. let x = pigs and y = chickens got it ?

OpenStudy (anonymous):

y

OpenStudy (anonymous):

in maths we use variables such as x and y , so that doin eqns will be easy or say we can even take for chickens = C and pigs = P if that is what you prefer :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

okay .. now chickens = C and pigs = P

OpenStudy (anonymous):

given that p + c = 169 (head) and 4p+2c = 398 (legs) because head is one for both pig and chicken and leg for a chicken is two and for a pig 4 .. got it ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay .. good now , Multiply 1st equation by 2 2 P + 2 C = 338 4 P + 2 C = 398 ( CHANGING SIGNS IN THE 2ND EQN) -2P - 0 = -60 2P = 60 P = 30 SO MUCH UNDERSTOOD ?

OpenStudy (anonymous):

@PJohnson89 dealing with a problem ?

OpenStudy (anonymous):

why does the sign change in the second half

OpenStudy (anonymous):

it is like that .. that is a part of math .. like let me ask you what is 2 * 3 ?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

yes .. just like that in quadratic eqns sign changes in the second eqn which is universally acceptd .. got it ?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

can you proceed now ? we have found that pig = 30 now you try doing for chickens ;)

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

do for chicken .. or you want me to do that too ?

OpenStudy (anonymous):

I will try it tomorrow I have to go to work in the morning

OpenStudy (anonymous):

hm..